A gas has a pressure of 0.370 atm at 50.0 °C. What is the pressure at standard temperature? (round answer to three sig figs)


1
Expert's answer
2021-08-18T17:12:04-0400

P1/T1=P2/T2

The standard temperature is 273.15 K.

T1 = 50.0 + 273.15 = 323.15 K

0.370/323.15 = P2/273.15

P2 = 0.313 atm